OpenStudy (mew55):

Maclaurin series (x+1)^4. i got the derivatives but having troubling setting it up :/

OpenStudy (anonymous):

i must be confused, because this is a polynomial. if you expand about 0 you get just what you would get if you multiply out

OpenStudy (mew55):

this is how my professor want us to write

OpenStudy (anonymous):

you can write \[1+4x+6x^2+4x^3+x^4\] but there are no terms after that a mclauren series for a polynomial is the polynomial itself

OpenStudy (mew55):

(x+1)^4 f(0) f(0)=1^4 f'(0)=4(x+1)^3 f''(0)=4*3(x+1)^2 f'''(0)=4*3*2(x+1) f''''(0)=4*3*2*1

OpenStudy (mew55):

oh now i understand *face palm* thank u XD

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

\(f'(x)=4(x+1)^3,f'(0)=4\)

OpenStudy (anonymous):

oh ok glad you got it. yw

OpenStudy (mew55):

XD